Circular No. 5940

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRB 940217 K. Hurley, University of California; D. Bertsch, T. Cline, B. Dingus,

J. Esposito, C. E. Fichtel, R. Hartman, S. Hunter, J. Mattox, C. von Montigny, R. Mukherjee, P. Sreekumar and D. Thompson, NASA/Goddard Space Flight Center; D. Kniffen, Hampden-Sydney College; J. Fierro, Y. C. Lin, P. Michelson, P. Nolan and T. Willis, Stanford University; E. Schneid, Grumman Aerospace Corporation; G. Kanbach, H. Mayer-Hasselwander and M. Sommer, Max Planck-Institut für Extraterrestrische Physik; G. Fishman, C. Kouveliotou and C. Meegan, NASA/Marshall Space Flight Center; and M. Boer and M. Niel, Centre d’Etude Spatiale des Rayonnements, Toulouse, report γ-ray observations by the EGRET instrument of the strong burst first detected by the BATSE experiment on Feb. 17.96021 UT, and on its annulus defined by timing using the ULYSSES/BATSE network. γ-rays above 30 MeV were recorded in the 170-s interval following the BATSE trigger. Additional events that appear to come from the burst direction were also recorded in the following 180 s. The preliminary position, based on only 15 photons, is α = 1^{h}56^{m}.8, δ = +3^{o}48′ (equinox 2000.0), with uncertainty radii of 0^{o}.3 and 0^{o}.7 for the 68-percent and 95percent confidence contours, respectively. The position is strongly influenced by the highest energy event at 4.4 GeV. The annulus defined by timing from ULYSSES and BATSE has a preliminary center at α = 21^{h}54^{m}.9, δ = +51^{o}57′ (equinox 2000.0), with a radius of 69^{o}.33 and a total width of 6′. The width of the annulus is expected to decrease with additional analysis. The nearest portion of the annulus is ∼ 0^{o}.1 from the center of the EGRET location. The EGRET location is also consistent with the position found by COMPTEL (Kippen et al., IAUC 5937) and by BATSE.

COMET McNAUGHT-RUSSELL (1993v) Recent total visual magnitude estimates indicate that this comet is

brighter than predicted on IAUC 5926: Feb. 13.46 UT, 9.5 (D. A. J. Seargent, The Entrance, N.S.W., 0.25-m reflector); 15.44, 10.4 (A. F. Jones, Nelson, New Zealand, 0.32-m reflector); 15.44, 9.1 (P. Camilleri, Cobram, Victoria, 20 × 80 binoculars); 16.39, 10.6 (Jones); 16.45, 9.0: (Camilleri).

PERIODIC COMET TEMPEL 1 (1993c) Total magnitude estimates: Feb. 6.15 UT, 15.1 (P. Pravec, Ondřejov,

Czech Republic, 0.65-m reflector + CCD); 10.45, 13.4 (A. Hale, near Orogrande, NM, 0.41-m reflector).

1994 February 24 Brian G. Marsden